Dillon Labs has asked its financial manager to measure the cost of each specific type of capital as well as the weighted average cost of capital. The weighted average cost is to be measured by using the following​ weights: 45​% ​long-term debt, 25​% preferred​ stock, and 30​% common stock equity​ (retained earnings, new common​ stock, or​ both). The​ firm's tax rate is 28​%.

Debt The firm can sell for ​$1030 a 16​-year, $1,000​-par-value bond paying annual interest at a 11.00​% coupon rate. A flotation cost of 4​% of the par value is required.

Preferred stock 8.00​% ​(annual dividend) preferred stock having a par value of ​$100 can be sold for ​$94.An additional fee of ​$2 per share must be paid to the underwriters.

Common stock The​ firm's common stock is currently selling for $59.43 per share. The stock has paid a dividend that has gradually increased for many​ years, rising from $2.25 ten years ago to the ​$3.84 dividend​ payment, D0​, that the company just recently made. If the company wants to issue new new common​ stock, it will sell them ​$3.50 below the current market price to attract​ investors, and the company will pay ​$4.00 per share in flotation costs.  

a- The after-tax cost of debt using the bond's yield to maturity (YTM) is ___% (Round to two decimal places.)
a1- The after tax-cost of debt using the approximation formula is ___% (Round to two decimal places.)

b- The cost of preferred stock is ___% (Round to two decimal places.)

c- The cost of retained earnings is ___% (Round to two decimal places.)
c1- The cost of new common stock is ___% (Round to two decimal places.)

d- Using the cost of retained earnings, the firm's WACC is ___% (Round to two decimal places.)
d1- Using the cost of new common stock, the firm's WACC is ___% (Round to two decimal places.)

PLEASE, Could you answer A, A1, B, C, C1, D ,D1 clearly like A= x% , B= x% , thank you.
Answer #1

A). Cost of debt using YTM:

FV (par value) = 1,000; PV (net price for new bond issue) = market price - (4% of par value) = 1,030 - (4%*1,000) = 990; N = 16; PMT (annual coupon) = annual coupon rate*par value = 11%*1,000 = 110, solve for RATE.

YTM = 11.14%

After-tax YTM = 11.14%*(1-Tax rate) = 11.14%*(1-28%) = 8.02%

A = 8.02% (Answer)

A1). Cost of debt using approximation formula:

Yield of bond = (annual interest + (1,000 - net bond price)/N)/((1,000 + net bond price)/2)

= (110 + (1,000-990)/16)/((1,000+990)/2) = 11.12%

After-tax yield = 11.12%*(1-28%) = 8.01%

A1 = 8.01% (Answer)

B). Cost of preferred stock = annual dividend/net price = (dividend rate*par value)/(price - flotation cost) = (8%*100)/(94-2) = 8.70%

B = 8.70% (Answer)

C). Cost of retained earnings:

D0 = 3.84; Price (P0) = 59.43

Dividend grows from 2.25 to 3.84 in 10 years. Let annual growth rate be g. Then,

[(1+g)^(1/10)] -1 = (3.84/2.25) - 1

Solving for g, we get g = 5.49%

Cost of retained earnings = (D0*(1+g)/P0) + g = (3.84*(1+5.49%)/59.43) + 6.119% = 12.31%

C = 12.31% (Answer)

C1). Cost of new equity:

D0 = 3.84; P0 = 59.43; Net price (Np) = P0 - 3.50 = 59.43-3.50 = 55.93; flotation cost (f) = 4; g = 5.49%

Cost of new equity = [D0*(1+g)/(Np-f)] + g = [3.84*(1+5.49%)/(55.93-4)] + 5.49% = 13.29%

C1 = 13.29% (Answer)

D). WACC = sum of weighted costs of capital = (45%*after-tax cost of debt) + (25%*cost of preferred stock) + (30%*cost of equity)

WACC (using retained earnings) = (45%*8.02%) + (25%*8.70%) + (30%*12.31%) = 9.47%

D = 9.47% (Answer)

D1). WACC (using cost of new common stock) = (45%*8.02%) + (25%*8.70%) + (30%*13.29%) = 9.77%

D1 = 9.77% (Answer)